Isooctane burns in air to give water and carbon dioxide (∆rH
0 = -10 992kJ/mol-rxn). What is the
enthalpy change if you burn 1.00 L of isooctane?
1
Expert's answer
2020-11-09T14:02:23-0500

The balanced reaction equation of isooctane combustion is:

2C8H18+27O2"\rightarrow" 16CO2+18H2O.

Therefore, the enthalpy in kJ/mol-rxn refers to 2 molecules of iso-octane and can be written as: -10 992 kJ/(2 mol C8H18).


The density "d" of isooctane is 690 g/L. To find the mass of isooctane, the following formulae can be used:

"m = d·V = 690\text{ g/L}·1.00\text{ L} = 690" g.


In order to convert the mass to the number of the moles, we must divide the mass by the molar mass of iso-octane: 114.23 g/mol.

"n = \frac{m}{M} =\frac{690 \text{ g}}{ 114.23\text{ g/mol}} = 6.04" mol.


Finally, the enthalpy change when 6.04 mol of isooctane is:

"∆H = -10992\text{ kJ}·\frac{6.04\text{ mol}}{2\text{ mol }C_8H_{18}} = -33198" kJ.

Answer: if you burn 1.00 L of isooctane, the enthalpy change is -33198 kJ.